OpenStudy (anonymous):

A poker hand (5 cards) is dealt. Find the chance that the cards are all hearts.

OpenStudy (kropot72):

There are 13 cards in the suit of hearts in a standard 52 card pack. This is sampling without replacement. P(5 hearts) = 13C5 / 52C5

OpenStudy (anonymous):

the answer is without replacement isnt it?

OpenStudy (kropot72):

\[P(5\ hearts)=\frac{13\times 12\times 11\times 10\times 9}{52\times 51\times 50\times 49\times 48}\]

OpenStudy (anonymous):

that's right thank you :)

OpenStudy (kropot72):

You're welcome :)